OpenStudy (anonymous):

A body of mass 5 kg is acted upon by two perpendicular forces 8 N and 6 N. Give the magnitude and direction of the acceleration of the body.

OpenStudy (anonymous):

@Vaidehi09 @DLS

OpenStudy (dls):

(Net force)^2 = 6^2 + 8^2 = 10^2 net force = 10 N

OpenStudy (anonymous):